Greg 'acquired' $25,000 and invested the money into two seprate accounts. In one of the accounts he earned 5.5% and in the second he earned 9%. In a year he took a total of $26,842.25 out of the accounts, closing them out. How much money did he invest at each rate?
Answer by stanbon(75887)

Greg 'acquired' $25,000 and invested the money into two seprate accounts. In one of the accounts he earned 5.5% and in the second he earned 9%. In a year he took a total of $26,842.25 out of the accounts, closing them out. How much money did he invest at each rate?
---------------------------------------
Equation:
interest + interest = interest
0.055x + 0.09(25000-x) = 26842.25-25000
Multiply thru by 1000 to get:
55x + 90*25000 - 90x = 1842250
-35x = -407750
x = $11,650 (amt invested at 5.5%
25000-11,650 = $13,350 (amt. inveted at 9%)
=============================================
cheers,
Stan H.